OpenStudy (anonymous):

Solve: 5(x − 1) = 3(x + 2) − 7

OpenStudy (compassionate):

x = 2

OpenStudy (compassionate):

5(x − 1) = 3(x + 2) − 7 5x - 5 = 3x + 6 - 7 7 - 6 - 5 = 3x - 5x -4 = -2x -- -- x = 2 -4 -4